hi, i need some help with these equations and if possible express the answer as on ordered pair....
x-y=4
2x+y=5
thanks in advance for the help!!

Answer by stanbon(75887)

1st: x-y=4
2nd: 2x+y=5
----------
Add 1st and 2nd to get:
3rd: 3x=9
x=3
Substitute into 1st or 2nd to solve for y, as follows:
3-y=4
y=-1
Solution: (3,-1)
Cheers,
Stan H.